THREE+THREE +FOUR=ELEVEN What numbers could take the place of the letters to make the equation work?
Answer by Edwin McCravy(20086)



There are four different solutions:

THREE + THREE + FOUR = ELEVEN
46511 + 46511 + 8295 = 101317

THREE + THREE + FOUR = ELEVEN
74611 + 74611 + 2096 = 151318

THREE + THREE + FOUR = ELEVEN
78011 + 78011 + 5390 = 161412

THREE + THREE + FOUR = ELEVEN
84011 + 84011 + 3590 = 171612

Edwin